# Imprints of quantum gravity effects on gravitational waves: a comparative study using extreme mass-ratio inspirals

### v1

@@ III.2 Fluxes and Evolution
Figure 5: Time evolution of the dephasing $\Delta\Phi={\Phi_{\phi}}_{\mathrm{I,II}}-{\Phi_{\phi}}_{\mathrm{Sch}}$ for various initial eccentricities at a fixed quantum gravity value of $\zeta=1/100$, where the inset shows an enlarged view around one year of orbital evolution. The left panel corresponds to the LQG-BH I background, and the right panel to the LQG-BH II background. The initial parameters $\left\{M,m,p_{0},{\Phi_{\phi}}_{0},{\Phi_{r}}_{0}\right\}$ are set to $\left\{10^{6}M_{\odot},10M_{\odot},10,0,0\right\}$ correspondingly.
Table 1: Summary of the accumulated dephasing $\Delta\Phi={\Phi_{\phi}}_{\mathrm{I,II}}-{\Phi_{\phi}}_{\mathrm{Sch}}$ over a one-year evolution, with fixed initial conditions $\left\{M,m,p_{0},{\Phi_{\phi}}_{0},{\Phi_{r}}_{0}\right\}=\left\{10^{6}M_{\odot},10M_{\odot},10,0,0\right\}$, for different initial eccentricities $e_{0}$ and quantum gravity parameters $\zeta$.
@@
Thus far, the influence of $\zeta$ on the waveform phases has been assessed through relative dephasing. In the following section, we proceed to synthesize and study the full time-domain EMRI gravitational-wave signals applicable to detection scenarios.
@@ IV.1 Set up
Using the numerical framework established above, we generate the $h^{+}$ polarization mode of EMRI waveforms for the Schwarzschild background and for LQG-BH backgrounds with various values of $\zeta$, assuming an initial eccentricity of 0.1 and a mass ratio of $10^{-5}$. Apart from these parameters, the remaining initial parameters required for waveform generation are summarized in Table 2. The resulting signals at the initial moment and after one year of evolution are presented in Fig. 6, where solid colored curves represent the quantum-corrected cases and dashed curves indicate the Schwarzschild reference. The left panel reveals that at the initial time, the waveforms in both BH-I and BH-II geometries are nearly indistinguishable from the Schwarzschild result, regardless of $\zeta$ values. After one year of evolution, the BH-I background with larger $\zeta$ exhibits clear phase separation from the Schwarzschild case, whereas the deviation remains negligible for $\zeta=1/1000$ (upper-right panel). By contrast, in the BH-II scenario, even comparatively large $\zeta$ values produce only minimal departures (bottom-left panel). These observational trends are consistent with the phase-difference analysis detailed in Sec. III.2.
These findings further suggest that cumulative phase evolution in the BH-I background enhances sensitivity to possible quantum gravity effects. In other words, GW observations of EMRIs in such spacetimes could produce signals distinct from those in the Schwarzschild case, whereas in the BH-II background, although the parameter $\zeta$ influences the waveforms, its impact is weak and likely beyond observational detectability.
Figure 6: The $h^{+}$ polarization of EMRI waveforms for different values of the quantum gravity parameter $\zeta$, computed under the initial conditions listed in Table 2. Results for the BH-I and BH-II spacetimes are shown in the upper and lower panels, respectively. The left column illustrates the initial signals, whereas the right column presents those after one year of evolution. Dashed curves indicate the corresponding Schwarzschild-case EMRIs.
@@ IV.2 Data analysis with faithfulness
On the right-hand side of the above expression, the coefficient ${\sqrt{3}}/{2}$ is a scale factor arising from the $60^{\circ}$ opening angle between LISA’s arms. The waveform components $h_{n}^{+}\left(t\right)$ and $h_{n}^{\times}\left(t\right)$ are already defined in Eqs. (57) and (58), ${F_{\rm{d1},\rm{d2}}^{+}}\left(t\right)$ and ${F_{\rm{d1},\rm{d2}}^{\times}}\left(t\right)$ are the interferometer beam-pattern functions, which depend on the trigonometric combinations of the source location $\left\{\theta_{\rm{S}},\phi_{\rm{S}}\right\}$ and the orbital angular momentum direction $\left\{\theta_{\rm{L}},\phi_{\rm{L}}\right\}$ in the ecliptic coordinate system, with explicit expressions provided in Refs. Barack:2003fp ; Cutler:1997ta ; Apostolatos:1994mx . It is important to note that, since this work is restricted to non-rotating BH spacetimes, the precession of $\hat{\mathbf{L}}\left(t\right)$ induced by the central BH’s spin is neglected, making it appropriate to treat $\left\{\theta_{\rm{L}},\phi_{\rm{L}}\right\}$ as fixed parameters.
Furthermore, in modeling the LISA gravitational-wave response, we incorporate the Doppler phase modulation generated by the detectors’ annual motion around the Sun, which becomes increasingly relevant for long observation times. This modulation is introduced in Eq. (60) by the replacement: $n\Phi_{r}\left(t\right)\to n\left(\Phi_{r}\left(t\right)+\Phi_{\rm{D}}\left(t\right)\right)$, where the Doppler phase is defined as Cutler:1997ta ; Barack:2003fp
@@
Figure 7: Faithfulness $\mathcal{F}$ between the EMRI waveforms generated in the LQG and Schwarzschild BH backgrounds, shown as functions of the parameter $\zeta$ for initial eccentricities $e_{0}=0.01$ and $e_{0}=0.1$. The waveform duration corresponds to one year of evolution from the onset of the inspiral. The black dashed line denotes the distinguishability threshold $\mathcal{F}_{\mathrm{thr}}=0.996$, associated with SNR of 30. The inset provides an enlarged view of the curves near this threshold.
Table 3: The lower limit values of parameter $\zeta$ for LISA discriminating quantum gravity effect under different eccentricity initial values, with the observational threshold of $\mathcal{F}$ is 0.996.
 Type-1 BH   $e_{0}\quad$ $\zeta\quad$
@@
 Type-2 BH   $e_{0}\quad$ $\zeta\quad$
@@ Acknowledgements.
R.T. Chen is especially grateful to Yunlong Liu for valuable discussions. This work is supported by National Key R&\&D Program of China (No. 2023YFC2206703), the Natural Science Foundation of China under Grant Nos. 12375055, 12505078 and 12505085, the Jiangsu Postgraduate Research and Practice Innovation Program under Grant No. KYCX25-3925, the China Postdoctoral Science Foundation (No. 2025T180931), and the Jiangsu Funding Program for Excellent Postdoctoral Talent (No. 2025ZB705).
@@ Appendix A Eccentricity-Dependent Dephasing in LQG-BH II
In section III.2, we have investigated the relative dephasing in LQG-BH spacetimes during the inspiral evolution within EMRI. From the results of LQG-BH II, we find an interesting behavior of the waveform phase when varying the initial eccentricity while keeping the parameter $\zeta=1/100$ fixed.
Here, we further demonstrate this dephasing behavior for a range of initial eccentricities. It is evident that for small eccentricities, quantum gravity corrections lead to a phase advance relative to the Schwarzschild waveform during the orbital evolution. As the eccentricity increases, this trend reverses, and the phase shift becomes a delay that grows with larger $e_{0}$. The transition is particularly evident for approximately $e_{0}=0.08$ and above. As discussed in Refs. Bonga:2019ycj ; Barsanti:2022vvl , the phase difference $\Delta\Phi\gtrsim 0.1$ radian is considered resolvable for a GW signal with $\mathrm{SNR}\approx 30$, which implies that such characteristics may serve as distinctive signatures in waveform templates for GW observations sensitive to quantum gravity effects for larger $\zeta$ in BH-II.

### v2

@@ III.2 Fluxes and evolution
Figure 5: Time evolution of the dephasing $\Delta\Phi={\Phi_{\phi}}_{\mathrm{I,II}}-{\Phi_{\phi}}_{\mathrm{Sch}}$ for various initial eccentricities at a fixed quantum gravity value of $\zeta=1/100$, where the inset shows an enlarged view around one year of orbital evolution.
Table 1: Accumulated EMRI dephasing $\Delta\Phi$ over one year evolution for different initial eccentricities $e_{0}$ and quantum gravity parameters $\zeta$.
@@
Thus far, the influence of $\zeta$ on the waveform phases has been assessed through relative dephasing. In the following section, we synthesize and study the full time-domain EMRI signals applicable to detection scenarios.
@@ IV.1 Set up
Using the numerical framework established above, we generate the $h^{+}$ polarization mode of EMRI waveforms for the Schwarzschild background and for LQG-BH backgrounds with various values of $\zeta$, assuming an initial eccentricity of 0.1 and a mass ratio of $10^{-5}$. Apart from these parameters, the remaining initial parameters required for waveform generation are summarized in Table 2. The resulting signals at the initial moment and after one year of evolution are presented in Fig. 6, where solid colored curves represent the quantum-corrected cases and dashed curves indicate the Schwarzschild reference. The left panel reveals that at the initial time, the waveforms in both BH-I and BH-II geometries are nearly indistinguishable from the Schwarzschild result, regardless of $\zeta$ values. After one year of evolution, the BH-I background with larger $\zeta$ exhibits clear phase separation from the Schwarzschild case, whereas the deviation remains negligible for $\zeta=1/1000$ (upper-right panel). By contrast, in the BH-II scenario, even comparatively large $\zeta$ values produce only minimal departures (bottom-right panel). These observational trends are consistent with the phase-difference analysis detailed in subsection III.2.
These findings further suggest that cumulative phase evolution in the BH-I background enhances sensitivity to possible quantum gravity effects. In other words, GW observations of EMRIs in such spacetimes could produce signals distinct from those in the Schwarzschild case, while in the BH-II background, although the parameter $\zeta$ influences the waveforms, its impact is weak and likely beyond observational detectability.
Figure 6: The $h^{+}$ polarization of EMRI waveforms for different values of the quantum gravity parameter $\zeta$.
@@ IV.2 Data analysis with faithfulness
On the right-hand side of the above expression, the coefficient ${\sqrt{3}}/{2}$ is a scale factor arising from the $60^{\circ}$ opening angle between LISA’s arms. The waveform components $h_{n}^{+}\left(t\right)$ and $h_{n}^{\times}\left(t\right)$ are already defined in Eqs. (57) and (58), ${F_{\rm{d1},\rm{d2}}^{+}}\left(t\right)$ and ${F_{\rm{d1},\rm{d2}}^{\times}}\left(t\right)$ are the interferometer beam-pattern functions, which depend on the trigonometric combinations of the source location $\left\{\theta_{\rm{S}},\phi_{\rm{S}}\right\}$ and the orbital angular momentum direction $\left\{\theta_{\rm{L}},\phi_{\rm{L}}\right\}$ in the ecliptic coordinate system Barack:2003fp ; Cutler:1997ta ; Apostolatos:1994mx . It is important to note that, since this work is restricted to non-rotating BH spacetimes, the precession of $\hat{\mathbf{L}}\left(t\right)$ induced by the central BH’s spin is neglected, making it appropriate to treat $\left\{\theta_{\rm{L}},\phi_{\rm{L}}\right\}$ as fixed parameters.
Furthermore, in modeling the GW response of LISA, we incorporate the Doppler phase modulation generated by the detectors’ annual motion around the Sun, which becomes increasingly relevant for long observation times. This modulation is introduced in Eq. (60) by the replacement: $n\Phi_{r}\left(t\right)\to n\left(\Phi_{r}\left(t\right)+\Phi_{\rm{D}}\left(t\right)\right)$, the Doppler phase $\Phi_{\rm{D}}\left(t\right)$ is defined as Cutler:1997ta ; Barack:2003fp
@@
Figure 7: Faithfulness $\mathcal{F}$ between EMRI waveforms in the LQG and Schwarzschild BH backgrounds. The inset shows an enlarged view near the threshold 0.996.
Table 3: Lower limits on the LQG parameter $\zeta$ detectable by LISA for different initial eccentricities $e_{0}$.
 Type-I BH   $e_{0}\quad$ $\zeta\quad$
@@
 Type-II BH   $e_{0}\quad$ $\zeta\quad$
@@ Acknowledgements.
R.T. Chen is grateful to Yunlong Liu and Prof. Andrea Maselli for their valuable discussions and clarification of related questions. This work is supported by National Key R&\&D Program of China (No. 2023YFC2206703), the Natural Science Foundation of China under Grant Nos. 12375055, 12505078 and 12505085, the Jiangsu Postgraduate Research and Practice Innovation Program under Grant No. KYCX25-3925, the China Postdoctoral Science Foundation (No. 2025T180931), and the Jiangsu Funding Program for Excellent Postdoctoral Talent (No. 2025ZB705).
@@ Appendix A Eccentricity-dependent dephasing in LQG-BH II
In subsection III.2, we have investigated the relative dephasing in LQG-BH spacetimes during the inspiral evolution within EMRI. From the results of LQG-BH II, we find an interesting behavior of the waveform phase when varying the initial eccentricity while keeping the parameter $\zeta=1/100$ fixed.
To illustrate this, we show in Fig. 8 the dephasing $\Delta\Phi$ for a range of initial eccentricities. It is evident that for small eccentricities, quantum gravity corrections lead to a phase advance relative to the Schwarzschild waveform during the orbital evolution. As the eccentricity increases, this trend reverses, and the phase shift becomes a delay that grows with larger $e_{0}$. The transition is particularly evident for approximately $e_{0}=0.08$ and above. As discussed in Refs. Bonga:2019ycj ; Barsanti:2022vvl , the phase difference $\Delta\Phi\gtrsim 0.1$ radian is considered resolvable for a GW signal with $\mathrm{SNR}\approx 30$, which implies that such characteristics may serve as distinctive signatures in waveform templates for GW observations sensitive to quantum gravity effects for larger $\zeta$ in BH-II.